OpenStudy (anonymous):

simplify 5 negative to squared

OpenStudy (abb0t):

\[5^{-2} ???\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

thats wat i ment

OpenStudy (abb0t):

\[5^{-2} = \frac{ 1 }{ 5^2 } = \frac{ 1 }{ 5 \times 5 } = ??\]